OpenStudy (anonymous):

The point P(15, 20) is on the terminal side of θ. Evaluate sin θ.

OpenStudy (anonymous):

help meeeeeeee

OpenStudy (anonymous):

Draw a triangle with a length of 15 and a height of 20. Then you can find your hypotenuse and use it to answer your question.

OpenStudy (anonymous):

so its 4/5?

OpenStudy (anonymous):

Hmmm...looks good.

OpenStudy (anonymous):

Good job!

OpenStudy (anonymous):

is it not 4/3?

OpenStudy (anonymous):

no, 4/3 is the tan.

OpenStudy (anonymous):

\[r = \sqrt{225+400}\]=25

OpenStudy (anonymous):

so 4/5 is the correct answer?

OpenStudy (anonymous):

yes, in reduced form, 20/25=4/5

OpenStudy (anonymous):

\[\sin \theta = 20\div25 = 0.8 = 4/5\] and so it seem to be correct